# Continuous Optimization for Feature Selection with Permutation-Invariant Embedding and Policy-Guided Search

### v1

@@ Abstract.
Feature selection removes redundant features to enhance both performance and computational efficiency in downstream tasks. Existing methods often struggle to capture complex feature interactions and adapt to diverse scenarios. Recent advances in this domain have incorporated generative intelligence to address these drawbacks by uncovering intricate relationships between features. However, two key limitations remain: 1) embedding feature subsets in a continuous space is challenging due to permutation sensitivity, as changes in feature order can introduce biases and weaken the embedding learning process; 2) gradient-based search in the embedding space assumes convexity, which is rarely guaranteed, leading to reduced search effectiveness and suboptimal subsets. To address these limitations, we propose a new framework that can: 1) preserve feature subset knowledge in a continuous embedding space while ensuring permutation invariance; 2) effectively explore the embedding space without relying on strong convex assumptions. For the first objective, we develop an encoder-decoder paradigm to preserve feature selection knowledge into a continuous embedding space. This paradigm captures feature interactions through pairwise relationships within the subset, removing the influence of feature order on the embedding. Moreover, an inducing point mechanism is introduced to accelerate pairwise relationship computations. For the second objective, we employ a policy-based reinforcement learning (RL) approach to guide the exploration of the embedding space. The RL agent effectively navigates the space by balancing multiple objectives. By prioritizing high-potential regions adaptively and eliminating the reliance on convexity assumptions, this search strategy effectively reduces the risk of converging to local optima. Finally, we conduct extensive experiments to demonstrate the effectiveness, efficiency, robustness and explicitness of our model. Our code and dataset are publicly accessible on GitHub111 https://anonymous.4open.science/r/CAPS-F01B.
isbn: 978-1-4503-XXXX-X/2018/06
@@ 3.3. Policy-Guided Multi-Objective Search
(11) $\mathbf{R}=\lambda(\mathcal{M}(X[\mathbf{f}^{+}])-\mathcal{M}(X[\mathbf{f}]))+(1-\lambda)\mathcal{N}[\mathbf{f}^{+}]),$
@@
where $\epsilon$ is a hyperparameter, $r_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})}{\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})}$ is the probability ratio of new policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})$ and old policy $\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})$ at $t$-th time step, $\mathbf{a}_{t}$ is the action of the actor network at $t$-th time step, $\hat{A}_{t}=V(\mathbf{s}_{t})-G_{t}$ is the estimated advantage function at $t$-th time step, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)$ is the clipping function to remove the unexpected incentive for moving $r_{t}(\theta)$ outside of the interval [1-$\epsilon$, 1+$\epsilon$]. Eventually, by taking the minimum of the clipped and unclipped objective, we ensure that $r_{t}(\theta)$ is always in a trust region. Once the actor network converges, we obtain the optimal policy $\pi^{*}$ that chooses the most appropriate action for the exploration. In other words, it can guide the embedding $\mathbf{E}$ in the learned space $\mathcal{E}$ to shift toward the direction resulting in high downstream ML task performance without any convexity assumption. From the training settings of aforementioned PPO agent, the well-trained agent $\mathcal{P}$ guarantees that the enhanced embedding $\mathbf{E}^{+}$ contributes to higher downstream ML task performance. Therefore, to identify the optimal feature indices, we should reconstruct the feature subset indices $\mathbf{f}^{+}$ from the enhanced embedding $\mathbf{E}^{+}$. In particular, we feed the $K$ enhanced embedding $[\mathbf{E}^{+}_{1},\mathbf{E}^{+}_{2},...,\mathbf{E}^{+}_{K}]$ into the well-trained decoder $\psi$ to reconstruct the enhanced feature subset indices $[\mathbf{f}^{+}_{1},\mathbf{f}^{+}_{2},...,\mathbf{f}^{+}_{K}]$. Consequently, we utilize the reconstructed feature subset indices to generate different feature subsets from the original dataset. Then, we input them into the downstream ML task to evaluate their performance and choose the feature subset indices with the highest performance as the optimal one $\mathbf{f}^{*}$.
@@ 4.1. Experimental Setup
We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point. (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning. Among the introduced baseline algorithms, K-Best and mRMR are categorized as filter methods. Lasso, RFE, and LassoNet are representative of embedded methods. GFS, SARLFS, and MARLFS are considered as wrapper methods. RRA, MCDM, and GAINS are classified as hybrid feature selection methods.
@@
4.2. Experimental Results
4.2.1. Overall performance.
4.2.2. Ablation Study.
@@ 4.2. Experimental Results
(a) Original Feature Set
(b) CAPS-selected Feature Subset
@@ 5. Related Works
Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time. Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005). However, such methods generally ignore the dependency and interactions among features. 2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Yang and Honavar, 1998; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996). However, such methods require excessive model training, resulting in high computational costs. 3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006). However, such methods heavily rely on the underlying model structure. Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a). For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset. However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space. To address these drawbacks, CAPS combines permutation-invariant embedding and policy-guided search. The permutation-invariant embedding reduces dependency on feature order, creating a more robust embedding space that preserves feature selection knowledge. The policy-guided search applies a reinforcement learning strategy to effectively explore this space, discovering better feature subsets without relying on strong convexity assumptions.

### v2

@@ Abstract.
Feature selection removes redundant features to enhance both performance and computational efficiency in downstream tasks. Existing methods often struggle to capture complex feature interactions and adapt to diverse scenarios. Recent advances in this domain have incorporated generative intelligence to address these drawbacks by uncovering intricate relationships between features. However, two key limitations remain: 1) embedding feature subsets in a continuous space is challenging due to permutation sensitivity, as changes in feature order can introduce biases and weaken the embedding learning process; 2) gradient-based search in the embedding space assumes convexity, which is rarely guaranteed, leading to reduced search effectiveness and suboptimal subsets. To address these limitations, we propose a new framework that can: 1) preserve feature subset knowledge in a continuous embedding space while ensuring permutation invariance; 2) effectively explore the embedding space without relying on strong convex assumptions. For the first objective, we develop an encoder-decoder paradigm to preserve feature selection knowledge into a continuous embedding space. This paradigm captures feature interactions through pairwise relationships within the subset, removing the influence of feature order on the embedding. Moreover, an inducing point mechanism is introduced to accelerate pairwise relationship computations. For the second objective, we employ a policy-based reinforcement learning (RL) approach to guide the exploration of the embedding space. The RL agent effectively navigates the space by balancing multiple objectives. By prioritizing high-potential regions adaptively and eliminating the reliance on convexity assumptions, this search strategy effectively reduces the risk of converging to local optima. Finally, we conduct extensive experiments to demonstrate the effectiveness, efficiency, robustness and explicitness of our model. Our code and dataset are publicly accessible on GitHub [^1] https://github.com/RayLiu1103/CAPS.
Automated Feature Selection; Representation Learning; Reinforcement Learning
journalyear: 2025copyright: ccconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.2 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3711896.3736891isbn: 979-8-4007-1454-2/2025/08ccs: Computing methodologies Feature selection
@@ 3.3. Policy-Guided Multi-Objective Search
(11) $\mathbf{R}=\lambda(\mathcal{M}(X[\mathbf{f}^{+}])-\mathcal{M}(X[\mathbf{f}]))+(1-\lambda)\mathcal{N}[\mathbf{f}^{+}],$
@@
where $\epsilon$ is a hyperparameter, $r_{t}(\theta)=\frac{\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})}{\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})}$ is the probability ratio of new policy $\pi_{\theta}(\mathbf{a}_{t}|\mathbf{s}_{t})$ and old policy $\pi_{\theta_{old}}(\mathbf{a}_{t}|\mathbf{s}_{t})$ at $t$-th time step, $\mathbf{a}_{t}$ is the action of the actor network at $t$-th time step, $\hat{A}_{t}=V(\mathbf{s}_{t})-G_{t}$ is the estimated advantage function at $t$-th time step, and $clip(r_{t}(\theta),1-\epsilon,1+\epsilon)$ is the clipping function to remove the unexpected incentive for moving $r_{t}(\theta)$ outside of the interval $\epsilon, 1$$\epsilon$]. Eventually, by taking the minimum of the clipped and unclipped objective, we ensure that $r_{t}(\theta)$ is always in a trust region. Once the actor network converges, we obtain the optimal policy $\pi^{*}$ that chooses the most appropriate action for the exploration. In other words, it can guide the embedding $\mathbf{E}$ in the learned space $\mathcal{E}$ to shift toward the direction resulting in high downstream ML task performance without any convexity assumption. From the training settings of aforementioned PPO agent, the well-trained agent $\mathcal{P}$ guarantees that the enhanced embedding $\mathbf{E}^{+}$ contributes to higher downstream ML task performance. Therefore, to identify the optimal feature indices, we should reconstruct the feature subset indices $\mathbf{f}^{+}$ from the enhanced embedding $\mathbf{E}^{+}$. In particular, we feed the $K$ enhanced embedding $\mathbf{E}^{+}_{1},\mathbf{E}^{+}_{2},...,\mathbf{E}^{+}_{K}$] into the well-trained decoder $\psi$ to reconstruct the enhanced feature subset indices $\mathbf{f}^{+}_{1},\mathbf{f}^{+}_{2},...,\mathbf{f}^{+}_{K}$]. Consequently, we utilize the reconstructed feature subset indices to generate different feature subsets from the original dataset. Then, we input them into the downstream ML task to evaluate their performance and choose the feature subset indices with the highest performance as the optimal one $\mathbf{f}^{*}$.
@@ 4.1. Experimental Setup
We compare our method with 12 widely-used feature selection methods: (1) K-Best (Yang and Pedersen, 1997) selects top-K features with the best feature scores; (2) mRMR (Peng et al., 2005) aims to select a feature subset that comprises of the features with greatest relevance and the least redundancy to the target; (3) LASSO (Tibshirani, 1996) conducts feature selection by regularizing model parameters, that is shrinking the coefficients of useless features into zero; (4) RFE (Granitto et al., 2006) selects features in a recursive fashion, that is removing the weakest features until the stipulated feature numbers is reached; (5) LASSONET (Lemhadri et al., 2021) conducts feature selection by training neural networks with novel objective function; (6) GFS (Leardi, 1996) conducts feature selection via genetic algorithms that is recursively generating a population based on a possible feature subset first, and then using a predictive model to assess it; (7) SARLFS (Liu et al., 2021a) utilizes an agent to select features among all features to alleviate computational costs and views feature redundancy and downstream task performance as incentives; (8) MARLFS (Liu et al., 2019) is an enhanced version of SARLFS, that is building a multi-agent system to select features with the same incentive settings with SARLFS, and each agent is associated with a single feature; (9) RRA (Seijo-Pardo et al., 2017) is a rank integration algorithm that first gathers several selected feature subsets, and then integrates them based on the their rank in terms of statistical information; (10) MCDM (Hashemi et al., 2022) first combines the ranks of features from different feature selection algorithms to form a decision matrix, and then assigns scores to features based on this matrix for the purpose of more comprehensively and accurately selecting the most valuable features through considering multiple evaluation criteria; (11) GAINS (Xiao et al., 2023c) first constructs a embedding vector space by jointly optimizing an encoder-evaluator-decoder model, and then uses the gradient from well-evaluator moves the embedding point toward high performance direction to search the optimal feature subset embedding point. (12) MEL (Wang et al., 2024) divides the parent population into two sub-populations that search for optimal feature subsets independently and uses the mutual influence and knowledge sharing between the sub-populations, integrating multi-task learning with evolutionary learning. Among the introduced baseline algorithms, K-Best and mRMR are categorized as filter methods. Lasso, RFE, and LassoNet are representative of embedded methods. GFS, SARLFS, and MARLFS are considered as wrapper methods. RRA, MCDM, and GAINS are classified as hybrid feature selection methods.
@@
4.1.4. Experimental Settings.
We conducted all experiments on the Windows 11 operating system, AMD Ryzen 5 5600X CPU, and NVIDIA GeForce RTX 3070Ti GPU, with the framework of Python 3.10.15 and PyTorch 2.5.1 (Paszke et al., 2019).
4.1.5. Hyperparameters and Reproducibility.
To collect potential feature subsets and their corresponding model performance, we run MARLFS for 300 epochs. To increase the diversity of training data, we augment the collected feature subset-accuracy records by permuting the order of each feature subset indices 25 times. The Encoder comprises of two layers of ISAB. The Decoder comprises of 3 components, which are PMA, MAB and rFF. The number of multi-attention head is 4. The embedding size of each feature index is 128. To train the encoder and decoder, we set batch size as 64, the step size as 0.001, and the dimension of seed vectors and inducing points as 32 respectively. During the search process, we used the top 25 feature selection records as the initial search seeds to search for the optimal embedding vector. To conduct stable and robust search, we set search epoch as 10, batch size as 512, learning rate of the actor as 0.0003, learning rate of the critic as 0.001, reward trade-off as 0.1, reward discounted factor as 0.99, search step as 1000, and the clipping ratio as 0.2 respectively.
@@
4.2. Experimental Results
4.2.1. Overall performance.
4.2.2. Ablation Study.
@@ 4.2. Experimental Results
(a) Clipping Trade-off
(b) Reward Trade-off
Figure 7. The hyperparameter sensitivity test on SpectF.
@@
4.2.7. Study of the Hyperparameter Sensitivity of Search Process.
We conduct this experiment to study the sensitivity of the search process performance to the the clipping trade-off $\epsilon$ and the reward trade-off $\lambda$. As part of the surrogate objective, the clipping trade-off $\epsilon$ can stabilize training and prevent excessively large updates to the original policy. Small value of $\epsilon$ results in stricter policy updates, ensuring that each step is small and stable, but at the risk of limiting exploration and making it harder for the policy to improve. Conversely, larger value of $\epsilon$ results in more generative policy updates, facilitating quicker exploration of the policy space. In other words, a smaller $\epsilon$ prioritizes exploitation over exploration, and vice versa. The reward trade-off $\lambda$ is designed for the RL agent to balance the attention on the feature subset performance and the total length of the subset from equation 11. In particular, a larger $\lambda$ prioritizes the feature subset performance over total length, and vice versa. We set both $\epsilon$ and $\lambda$ vary from 0.1 to 0.9, then train the CAPS on Spectf. The model performance is shown in Figure 7 in terms of Precision, Recall and F1-Score. Overall, we observed that CAPS achieves best performance when $\epsilon$ is set to 0.2 and $\lambda$ is set to 0.1. From an empirical perspective, PPO has been shown to perform robustly across a wide range of tasks, with $\epsilon$ typically set to 0.2 in many standard tasks (Schulman et al., 2017), which aligns with our findings. As for the reward trade-off $\lambda$, the model performance remains relatively stable across most values of $\lambda$ and occasionally increases when $\lambda$ is set to 0.1. These findings provide insights into how $\epsilon$ and $\lambda$ influence the search results and how we should choose optimal values for them.
@@ 5. Related Works
Feature selection screens discriminative features from high dimensional data, improving model efficiency and performance by reducing dimensionality and shortening training time. Conventional feature selection methods are categorized into three main types: 1) Filter methods evaluate feature importance using statistical metrics (e.g., correlation, mutual information and chi-square tests) (Yang and Pedersen, 1997; Hall, 1999; Forman et al., 2003; Yu and Liu, 2003; Peng et al., 2005; Biesiada and Duch, 2008; Ding et al., 2014). However, such methods generally ignore the dependency and interactions among features. 2) Wrapper methods formulate feature selection as a search problem, training a machine learning (ML) model as the evaluation criterion and selecting the subset that yields the best performance (Liu et al., 2021b; Kim et al., 2000; Narendra and Fukunaga, 1977; Kohavi and John, 1997; Leardi, 1996). However, such methods require excessive model training, resulting in high computational costs. 3) Embedded methods integrate feature selection into the model training process as a part of the optimization objective to evaluate the importance of features (Tibshirani, 1996; Sugumaran et al., 2007; Lemhadri et al., 2021; Granitto et al., 2006; Kumagai et al., 2022; Koyama et al., 2022). However, such methods heavily rely on the underlying model structure. Recently, generative methods have attracted substantial attention due to the success of generative AI (Xiao et al., 2023c; Ying et al., 2024a; Wang et al., 2022; Xiao et al., 2023b, a). For instance, (Xiao et al., 2023c) introduces a framework that encompasses a sequential encoder, an accuracy evaluator, a sequential decoder, and a gradient ascent optimizer, embedding discrete feature selection knowledge into a continuous embedding space and conducting gradient-ascent search process for the optimal feature subset. However, this approach is limited by: 1) incapable of encoding the permutation invariance into the embedding space; 2) the reliance on the convexity assumptions about the feature subset embedding space. To address these drawbacks, we integrate a permutation-invariant embedding mechanism with a policy-guided search strategy in CAPS. The permutation-invariant embedding is designed to mitigate the sensitivity to feature ordering, constructing a more robust embedding space that preserves feature selection knowledge. The policy-guided search leverages a reinforcement learning strategy to effectively explore this space, identifying better feature subsets without relying on strong convexity assumptions.
(a) Original Feature Set
(b) CAPS-selected Feature Subset